# On-sky demonstration of an ultra-fast intensity interferometry instrument utilising hybrid single photon counting detectors

### v1

@@ 3.1 Lab test
Before testing our setup, we measured the timing resolution of our detection system (TDC, HPDs and CFDs) in the laboratory using a pulsed fs-laser to be 42.442.442.4\,42.4ps, much lower than the timing jitter of typical detection systems (see Appendix  A). Together with the filter transmission spectra supplied by the filter manufacturers this measurement allows the calculation of both the expected coherence time and expected shape of the bunching peak using equations Eq 1, 2, 3, and 4. We calculate the expected coherence time to 0.2120.2120.212\,0.212ps, and plot the expected shape of the bunching peak in each of the measurement results shown below. A $2f-2f$ configuration test of our setup, as described in Sec. 2.4, confirmed our expectation and revealed a coherence time of $(0.21\pm 0.02)\,$ps (see Fig. 9).
@@ 3.2 Bunching of A-type stars
Once the setup was mounted on the Omicron telescope, we observed three bright A-type stars: Vega, Altair, and Deneb. The star light was measured starting the night of 2023-08-09 and ending the night of 2023-08-18. During one night no measurements could be taken (2023-08-16) due to bad weather conditions, and during one night the results were compromised by a malfunction of the quTAG (2023-08-11), leaving us with eight good measurement nights. Throughout all these nights the sky was very clear, with only some slight high layer clouds in rare instances. A summary of the measurements is given in Tab. 1, showing the apparent magnitudes, the coherence time as well as its error, the observation time, and SNR for each star. A significant portion of our observation time (12.112.112.1\,12.1h) was spent on Vega, aiming for bunching at a high SNR. In addition Vega was observed at the beginning of each night starting 2023-08-12 for roughly 30 minutes, except for 2023-08-18, using the large photon count rate to ensure our detection electronics were working correctly. The largest part of our observation time was spent on Altair (29.929.929.9\,29.9h), attempting to resolve bunching for a star with slightly larger apparent magnitude. We spent the final nights of our campaign observing Deneb, in order to observe bunching for an even dimmer star.
The bunching measurement results for each of the stars are plotted in Fig 5. These plots show the measurement result with error bars generated by measuring the root mean square error (RMSE) of the baseline of the correlation histogram far from the correlation peak. We further show the expected measurement result (blue crosses), as well as a Gaussian fit to the data which is only used for centering the bunching peak with respect to zero time delay. In all cases the measurement results fit to our expectation very well, deviating from the expectation by less than $1\,\sigma$. The coherence times listed in Tab. 1 and shown in the plot legends were obtained by numerical integration between the $3\,\sigma$ borders of the expected measurement result. This guarantees equal integration borders for all three measurements. The uncertainties of these coherence times were obtained via appropriate error propagation of the bin uncertainties shown as error bars. Due to the differing count rates owed to different apparent magnitudes and the varying observation times each bunching measurement shows a different SNR. We obtained the highest SNR for Vega of 12, and managed to observe bunching with an SNR of 7.3 and 2.7 for Altair and Deneb respectively.
@@
Despite introducing an appropriate cable delay and taking appropriate measures to shield our electronics from the RF background of the telescope dome, we observed some deviation from shot noise when evaluating the RMSE for each of our measurements. The observed RMSE over the number of counts per bin $N$ is shown in Fig. 6. Ideally all RMSE curves should show shot noise closely following a square root curve. For Vega this curve is displayed in the top part of Fig. 6. Plotting the residuals divided by their shot noise value in the lower part of Fig. 6 shows that the relative residuals are initially close to zero for all observed stars, but around $N=0.25\times 10^{6}$ the deviations grow. Furthermore, the relative residuals are larger for targets at higher count rates, with Vega deviating 6% from shot noise at the end of its observation time, Altair deviating 1.5%, and Deneb nearly not deviating at all. While this discourages observations at higher detected photon rates with our current time tagger, we also note that the relative residuals seem to be rather stable once settled in, still allowing high SNR bunching measurements. Similar deviations from shot noise have been measured in our laboratory test (cf. Appendix B). We thus conclude that this deviation from shot noise is most likely intrinsic to our detection electronics.
@@
Figure 6: Root mean square error (RMSE) (top) and re-scaled residuals of the RMSE from shot noise (bottom) plotted over the number of counts per bin $N$. The RMSE is expected to be shot noise $\mathrm{RMSE}=\sqrt{N}$ for $N$ counts per histogram bin. The upper plot shows deviations from shot noise in Vega’s curve for higher count rates. This is due to the limits of the detection system. These deviations are shown in greater detail for all observed stars in the lower plot. While the residuals are initially close to zero for all stars, the relative residuals are larger for targets with higher detected count rates. Even for the high SNR measurement recorded for Vega, the RMSE is at most 6% larger than the shot noise expectation.
@@
Star App. Magnitude meas. $\tau_{c}$ [ps] Obs. Time [h] SNR
@@ 3.3 Advantages of our setup
The short optical path of our setup, i.e., the distance from the entrance point of the light to the detector allows a straightforward alignment. The alignment is done by adjusting merely the telescope’s focus and its pointing/tracking of the star, in order for the beam to impinge centrally. All optical components were pre-aligned in the lab and no further adjustments of optics inside the setup were necessary.
@@
Tab. 2 shows the observed count rates for all stars evaluated over all measurement nights. The median, maximum, and minimum count rate are calculated in a standard fashion, taking all data into account uncorrected. Note that the rather small minima for the count rates are due to bad weather conditions, i.e. clouds covering the star for a short time during the night. The deviation in Tab. 2 is calculated after correcting for the deterministic change in count rate due to the star’s motion along the night sky. For all observed stars, the standard deviation of the count rate is orders of magnitude smaller than the median observed count rate, indicating fairly stable coupling of the star light to our detectors. For Vega and Deneb, we observe a larger fluctuation in count rate than for Altair, indicating slightly worse observation conditions. This is also reflected by the more pronounced loss in count rate when comparing minimum to median count rate for Vega and Deneb with respect to Altair.
@@
Figure 7: Count rate at the detectors over time during the night of the 13th of August. The count rate is averaged every 303030\,30s. Three different stars were observed. One can clearly see that each measurement was quite stable, i.e. tracking of the star was never lost. This is due to the stability of the equatorial yoke mount of the telescope and the rather large active area of the HPDs. The time-evolutions of the count rates for Altair and Deneb are mainly driven by the air mass evolution due to sidereal motion. For Vega, the flux variation is more likely to be due to atmospheric transmission fluctuations.
@@ 4 Conclusion and Outlook
We have successfully measured temporal photon correlations and photon bunching of three different bright A-type stars - Vega, Altair, and Deneb. In all cases the observed coherence time fits well to both the pre-calculated expectations as well as the value measured in preceding laboratory tests. Our setup features an impressively low timing resolution of 42.442.442.4\,42.4ps. The maximum SNR was 12, achieved for Vega after 12.112.112.1\,12.1h of observation. For Altair and Deneb, the SNR was significantly lower, i.e., 7.3 and 2.7, respectively. Taking into account the apparent magnitudes of the observed stars in the U and B bands, we obtained more than 79 per cent of the expectation for each target due to the efficient coupling of the starlight to our detector.
In that respect, our setup shows notable advantages with respect to fibre coupled detector configurations. These will also hold when measuring spatial photon correlations, i.e. when two setups are used at two telescopes. Due to the large active area of our photodetectors the entire setup can be easily aligned at the telescope without need for readjustments of the optical components. The large area also allows for stable coupling of the light field to the detectors, decreasing the sensitivity to both tracking errors and atmospheric seeing. Spatial correlation measurements are feasible on the instrumentation side and promising for testing common astronomical models for stars in the blue. Our setup worked well at the Omicron telescope, hence, we plan to perform spatial correlation measurements in the near future in the blue at C2PU.
For these measurements, we plan to use a new type of detector based on a Photonis Hi-QE photocathode and multichannel plate (MCP) amplification. These detectors combine the advantages of a large active area of 888\,8mm, a good timing resolution of $<75\,$ps, and a nearly deadtime free operation. The upgrade to these detectors would increase our SNR per unit measurement time by a factor of $\approx 3$ due to the Hi-QE photocathode’s enhanced quantum efficiency. The latter is 50 per cent greater compared to conventional Bialkali photocathodes, like the ones used in our HPDs. A further benefit is offered by the fact that such detectors are virtually deadtime free due to the ability of the large MCP to support multiple confined electron avalanches within a short timescale. This prevents the loss of potentially detected photons due to a previous detection, which is an advantageous feature even at the low event rates considered in this paper.
@@
Using Hi-QE instead of HPD detectors and observing a magnitude 00 A-type star like Vega, spatial correlations with SNR 5 (10) are expected to be carried out within 0.880.880.88\,0.88h (3.513.513.51\,3.51h) per baseline. This means that for a mag. 00 A-type star and two 111\,1m diameter telescopes, high resolution intensity interferometry measurements within one observation night are within reach. For a slightly dimmer star like Altair, the required observation time for the same SNR values increase to 6.526.526.52\,6.52h per baseline for an SNR of 5, which is still feasible given the combination of multiple observation nights. Measuring with two detectors per telescope will give access to four independent spatial correlation measurements with the same baseline, which can be added up to decrease the bin error by a factor of two over a single temporal correlation measurement. This will however not increase the SNR of the spatial photon correlation measurement, as the squared visibility for most bright A-type stars is about $0.5$ for the 151515\,15m on-ground baseline of C2PU, reducing the bunching peak height by that factor. In fact these two contributions cancel out, leading to the same SNR as in a temporal correlation measurement using a single telescope.
Acknowledgements
@@ Appendix B Lab test results
In terms of measured coherence time, the laboratory test fits our expectations extremely well, with $(0.21\pm 0.02)\,$ps and an expected coherence time of 0.210.210.21\,0.21ps (cf. Fig. 9). To measure a normalized second order photon correlation function, like the one shown in Fig. 9, some preprocessing is necessary. A raw correlation measurement is shown in Fig. 10. Note that here the second order correlation function $\Gamma^{(2)}(\tau_{\text{raw }})$ is not normalized. Besides the correlation peak (highlighted in green), some other regularly spaced peaks (highlighted in red) are visible. These peaks likely correspond to the frequency of the readout clock of the quTAG. The sections between these clock spikes are reasonably close to being shot noise limited. Thus, a 333\,3m cable delay is introduced such that the correlation peak lies inside one of these regions. The clock spikes shaded in red, as well as the region of the correlation peak shaded in green, were not used for the RMSE evaluations presented in this paper. Similarly, only the non-shaded ares of the raw correlation histogram were used in the determination of the average counts per bin, used to calculate the normalized second order correlation function.
Even in the laboratory, our detection electronics did not conform to shot noise perfectly. In fact, the deviations were more significant than at the telescope, finally saturating at a RMSE 15 per cent larger than the shot noise expectation (cf. Fig 11). The step in the relative RMSE at about $2\times 10^{6}$ counts per bin is likely due to the long uninterrupted uptime (more than 424242\,42h) of the detection system, as parameters like the count rate or the ambient temperature did not fluctuate during the lab test.
@@ Appendix C HPD quantum efficiency
The quantum efficiency measurement is performed in three steps: first a calibrated photodiode together with a pico-amperemeter is used to collect a reference flux measurement at the desired wavelengths. This photodiode has an area larger than the active area of the HPD. In step two the same photodiode is used to measure the attenuation of a set of ND filters at the same wavelengths. These ND filters are necessary to attenuate the XBOs flux to a level the HPD can tolerate. In step three the HPD is used in the light beam attenuated by the ND filters calibrated in step two. For this measurement the HPD is connected directly to the quTAG without any signal conditioning by a CFD. Instead the pulse detection threshold at the quTAG is set to be just above the voltage level of HPD noise, insuring all single photon detection signals are discretised. The data necessary for the quantum efficiency measurement is acquired continuously, with the monochromators optical shutter acting as the signal to changing of the wavelength incident on the HPD.

### v2

@@ 3.1 Lab test
Before testing our setup, we measured the timing resolution of our detection system (TDC, HPDs and CFDs) in the laboratory using a pulsed fs-laser to be 42.442.442.4\,42.4ps (see Appendix  A), which is low compared to typical detection systems that are on the order of 100ps to ns. Together with the filter transmission spectra supplied by the filter manufacturers this measurement allows the calculation of both the expected coherence time and expected shape of the bunching peak using equations Eq 1, 2, 3, and 4. We calculate the expected coherence time to 0.2120.2120.212\,0.212ps, and plot the expected shape of the bunching peak in each of the measurement results shown below. A $2f-2f$ configuration test of our setup, as described in Sec. 2.4, confirmed our expectation and revealed a coherence time of $(0.21\pm 0.02)\,$ps (see Fig. 9).
@@ 3.2 Bunching of A-type stars
Once the setup was mounted on the Omicron telescope, we observed three bright A-type stars: Vega, Altair, and Deneb. The star light was measured starting the night of 9 August 2023 and ending the night of 18 August 2023. During one night no measurements could be taken (16 August 2023) due to bad weather conditions, and during one night the results were compromised by a malfunction of the quTAG (11 August 2023), leaving us with eight good measurement nights. Throughout all these nights the sky was very clear, with only some slight high layer clouds in rare instances. A summary of the measurements is given in Table 1, showing the apparent magnitudes, the coherence time as well as its error, the observation time, and S/N for each star. A significant portion of our observation time (12.112.112.1\,12.1h) was spent on Vega, aiming for bunching at a high S/N. In addition Vega was observed at the beginning of each night starting 2023-08-12 for roughly 30 minutes, except for 2023-08-18, using the large photon count rate to ensure our detection electronics were working correctly. The largest part of our observation time was spent on Altair (29.929.929.9\,29.9h), attempting to resolve bunching for a star with slightly larger apparent magnitude. We spent the final nights of our campaign observing Deneb, in order to observe bunching for an even dimmer star.
The bunching measurement results for each of the stars are plotted in Fig 5. These plots show the measurement result with error bars generated by measuring the root mean square error (RMSE) of the baseline of the correlation histogram far from the correlation peak. We further show the expected measurement result (blue crosses), as well as a Gaussian fit to the data which is only used for centering the bunching peak with respect to zero time delay. In all cases the measurement results fit to our expectation very well, deviating from the expectation by less than $1\,\sigma$. The coherence times listed in Table 1 and shown in the plot legends were obtained by numerical integration between the $3\,\sigma$ borders of the expected measurement result. This guarantees equal integration borders for all three measurements. The uncertainties of these coherence times were obtained via appropriate error propagation of the bin uncertainties shown as error bars. Due to the differing count rates owed to different apparent magnitudes and the varying observation times each bunching measurement shows a different S/N. We obtained the highest S/N for Vega of 12, and managed to observe bunching with an S/N of 7.3 and 2.7 for Altair and Deneb respectively.
@@
Despite introducing an appropriate cable delay and taking appropriate measures to shield our electronics from the RF background of the telescope dome, we observed some deviation from shot noise when evaluating the RMSE for each of our measurements. The observed RMSE over the number of counts per bin $N$ is shown in Fig. 6. Ideally all RMSE curves should show shot noise closely following a square root curve. For Vega this curve is displayed in the top part of Fig. 6. Plotting the residuals divided by their shot noise value in the lower part of Fig. 6 shows that the relative residuals are initially close to zero for all observed stars, but around $N=0.25\times 10^{6}$ the deviations grow. Furthermore, the relative residuals are larger for targets at higher count rates, with Vega deviating 6% from shot noise at the end of its observation time, Altair deviating 1.5%, and Deneb nearly not deviating at all. While this discourages observations at higher detected photon rates with our current time tagger, we also note that the relative residuals seem to be rather stable once settled in, still allowing high S/N bunching measurements. Similar deviations from shot noise have been measured in our laboratory test (cf. Appendix B). We thus conclude that this deviation from shot noise is most likely intrinsic to our detection electronics.
@@
Figure 6: Root mean square error (RMSE) (top) and re-scaled residuals of the RMSE from shot noise (bottom) plotted over the number of counts per bin $N$. The RMSE is expected to be shot noise $\mathrm{RMSE}=\sqrt{N}$ for $N$ counts per histogram bin. The upper plot shows deviations from shot noise in Vega’s curve for higher count rates. This is due to the limits of the detection system. These deviations are shown in greater detail for all observed stars in the lower plot. While the residuals are initially close to zero for all stars, the relative residuals are larger for targets with higher detected count rates. Even for the high S/N measurement recorded for Vega, the RMSE is at most 6% larger than the shot noise expectation.
@@
Star App. Magnitude meas. $\tau_{c}$ [ps] Obs. Time [h] S/N
@@ 3.3 Advantages of our setup
The short optical path of our setup, that is the distance from the entrance point of the light to the detector allows a straightforward alignment. The alignment is done by adjusting merely the telescope’s focus and its pointing/tracking of the star, in order for the beam to impinge centrally. All optical components were pre-aligned in the lab and no further adjustments of optics inside the setup were necessary.
@@
Table 2 shows the observed count rates for all stars evaluated over all measurement nights. The median, maximum, and minimum count rate are calculated in a standard fashion, taking all data into account uncorrected. Note that the rather small minima for the count rates are due to bad weather conditions, these are. clouds covering the star for a short time during the night. The deviation in Table 2 is calculated after correcting for the deterministic change in count rate due to the star’s motion along the night sky. For all observed stars, the standard deviation of the count rate is orders of magnitude smaller than the median observed count rate, indicating fairly stable coupling of the star light to our detectors. For Vega and Deneb, we observe a larger fluctuation in count rate than for Altair, indicating slightly worse observation conditions. This is also reflected by the more pronounced loss in count rate when comparing minimum to median count rate for Vega and Deneb with respect to Altair.
@@
Figure 7: Count rate at the detectors over time during the night of the 13th of August. The count rate is averaged every 303030\,30s. Three different stars were observed. One can clearly see that each measurement was quite stable, meaning. tracking of the star was never lost. This is due to the stability of the equatorial yoke mount of the telescope and the rather large active area of the HPDs. The time-evolutions of the count rates for Altair and Deneb are mainly driven by the air mass evolution due to sidereal motion. For Vega, the flux variation is more likely to be due to atmospheric transmission fluctuations.
@@ 4 Conclusion and outlook
We have successfully developed an ultra-fast intensity interferometry instrument using off the shelf components with a high timing resolution of 42.442.442.4\,42.4ps. We measured temporal photon correlations and photon bunching of three different bright A-type stars - Vega, Altair, and Deneb. In all cases the observed coherence time fits well to both the pre-calculated expectations as well as the value measured in preceding laboratory tests. The maximum S/N was 12, achieved for Vega after 12.112.112.1\,12.1h of observation. For Altair and Deneb, the S/N was significantly lower, that is 7.3 and 2.7, respectively. Taking into account the apparent magnitudes of the observed stars in the U and B bands, we obtained more than 79 per cent of the expectation for each target due to the efficient coupling of the starlight to our detector.
In that respect, our setup shows notable advantages with respect to fiber coupled detector configurations. These will also hold when measuring spatial photon correlations, that is when two setups are used at two telescopes. Due to the large active area of our photodetectors the entire setup can be easily aligned at the telescope without need for readjustments of the optical components. The large area also allows for stable coupling of the light field to the detectors, decreasing the sensitivity to both tracking errors and atmospheric seeing resulting in a nearly autonomous operation at the telescope. Spatial correlation measurements are feasible on the instrumentation side and promising for testing common astronomical models for stars in the blue. Our setup worked well at the Omicron telescope, hence, we plan to perform spatial correlation measurements in the near future in the blue at C2PU.
For these measurements, we plan to use a new type of detector based on a Photonis Hi-QE photocathode and multichannel plate (MCP) amplification. These detectors combine the advantages of a large active area of 888\,8mm, a good timing resolution of $<50\,$ps, and a nearly deadtime free operation. The upgrade to these detectors would increase our S/N per unit measurement time by a factor of $\approx 3$ due to the Hi-QE photocathode’s enhanced quantum efficiency. The latter is 50 per cent greater compared to conventional Bialkali photocathodes, such as the ones used in our HPDs. A further benefit is offered by the fact that such detectors are virtually deadtime free due to the ability of the large MCP to support multiple confined electron avalanches within a short timescale. This prevents the loss of potentially detected photons due to a previous detection, which is an advantageous feature even at the low event rates considered in this paper.
@@
Using Hi-QE instead of HPD detectors and observing a magnitude 00 A-type star such as Vega, spatial correlations with S/N 5 (10) are expected to be carried out within 0.90.90.9\,0.9h (3.53.53.5\,3.5h) per baseline. This means that for a mag. 00 A-type star and two 111\,1m diameter telescopes, high resolution intensity interferometry measurements within one observation night are within reach. For a slightly dimmer star such as Altair, the required observation time for the same S/N values increase to 6.56.56.5\,6.5h per baseline for an S/N of 5, which is still feasible given the combination of multiple observation nights. Measuring with two detectors per telescope will give access to four independent spatial correlation measurements with the same baseline, which can be added up to decrease the bin error by a factor of two over a single temporal correlation measurement. This will however not increase the S/N of the spatial photon correlation measurement, as the squared visibility for most bright A-type stars is about $0.5$ for the 151515\,15m on-ground baseline of C2PU, reducing the bunching peak height by that factor. In fact these two contributions cancel out, leading to the same S/N as in a temporal correlation measurement using a single telescope.
@@ Appendix B Lab test results
In terms of measured coherence time, the laboratory test fits our expectations extremely well, with $(0.21\pm 0.02)\,$ps and an expected coherence time of 0.210.210.21\,0.21ps (cf. Fig. 9). To measure a normalized second order photon correlation function, such as the one shown in Fig. 9, some preprocessing is necessary. A raw correlation measurement is shown in Fig. 10. Note that here the second order correlation function $\Gamma^{(2)}(\tau_{\text{raw }})$ is not normalized. Besides the correlation peak (highlighted in green), some other regularly spaced peaks (highlighted in red) are visible. These peaks likely correspond to the frequency of the readout clock of the quTAG. The sections between these clock spikes are reasonably close to being shot noise limited. Thus, a 333\,3m cable delay is introduced such that the correlation peak lies inside one of these regions. The clock spikes shaded in red, as well as the region of the correlation peak shaded in green, were not used for the RMSE evaluations presented in this paper. Similarly, only the non-shaded ares of the raw correlation histogram were used in the determination of the average counts per bin, used to calculate the normalized second order correlation function.
Even in the laboratory, our detection electronics did not conform to shot noise perfectly. In fact, the deviations were more significant than at the telescope, finally saturating at a RMSE 15 per cent larger than the shot noise expectation (cf. Fig 11). The step in the relative RMSE at about $2\times 10^{6}$ counts per bin is likely due to the long uninterrupted uptime (more than 424242\,42h) of the detection system, as parameters such as the count rate or the ambient temperature did not fluctuate during the lab test.
@@ Appendix C HPD quantum efficiency
The quantum efficiency measurement is performed in three steps: first a calibrated photodiode together with a pico-amperemeter is used to collect a reference flux measurement at the desired wavelengths. This photodiode has an area larger than the active area of the HPD. In step two the same photodiode is used to measure the attenuation of a set of ND filters at the same wavelengths. These ND filters are necessary to attenuate the XBOs flux to a level the HPD can tolerate. In step three the HPD is used in the light beam attenuated by the ND filters calibrated in step two. For this measurement the HPD is connected directly to the quTAG without any signal conditioning by a CFD. Instead the pulse detection threshold at the quTAG is set to be just above the voltage level of HPD noise, insuring all single photon detection signals are discretized. The data necessary for the quantum efficiency measurement is acquired continuously, with the monochromators optical shutter acting as the signal to changing of the wavelength incident on the HPD.